3. An incandescent lightbulb is rated at 100 Watts when plugged into a 110 V-rms household outlet. Calculate the resistance of the filament and the rms current. a. 12.22, 0.9A (rms) b. 102, 1A (rms) c. 1102, 1A (rms) d. 122Ω, 0.9 A (rms) e. 119, 1.1 A (rms)
Answer #1

power, P = V2/R

thus R = V2/P = 1102/100 = 121 ohms

rms current, I = V/R = 110/121 = 0.9A

so option d is correct